# Better Call SAL: Towards Learning to Segment Anything in Lidar

### v1

@@ 4.4 Lidar Panoptic Segmentation Evaluation
Finally, we compare SAL with state-of-the-art LPS methods on standard benchmarks [6, 18]. We evaluate SAL model in two settings: (i) zero-shot, utilizing the dataset vocabulary provided by [6, 18] as text prompts (detailed in Sec.˜C.1), and (ii) linear probing, where we train a (query) linear classification head directly on the respective dataset classes (while keeping backbone/decoder weights frozen). Importantly, zero-shot results are obtained without training on any labeled data. In contrast, all baselines are trained using human-labeled data, and are constrained to their respective, pre-defined class vocabularies.
@@
SemanticKITTI.  We report results in Tab.˜5 and focus this discussion on a comparison between fully-supervised, linearly probed, and zero-shot SAL models. When evaluating with the default class vocabulary, our zero-shot model reaches $23.6$ PQ or $40\%$ of the fully-supervised model ($59.5$ PQ). Linear probing yields $33.1$ PQ ($56\%$ of the fully-supervised model). As discussed in Sec.˜4.2, our method is mainly limited by the quality of the generated CLIP features. In practice, we observe that our zero-shot semantic classifier mainly confuses objects within sets of super classes, e.g., car vs. other-vehicle. Therefore, without any retraining, we prompt our zero-shot model with super classes. In this setting, our model reaches $46.3$ PQ, $57\%$ of the supervised model ($81.7$ PQ), and further improves to $55.3$ PQ via linear probing ($68\%$ of the supervised model).
@@
nuScenes.  We report our result in Tab.˜5 and reach similar conclusions. SAL reaches $31.3$ PQ on default classes, $44\%$ of the supervised model ($70.5$ PQ). With linear probing, we reach $40.1$ PQ ($57\%$ of fully supervised). As with SemanticKITTI, results consistently improve when evaluated with super-classes. In this setting, zero-shot and linear probing yields $42.1$ and $59.2$ PQ, respectively. Note that SAL when supervised with labeled data, significantly outperforms MaskPLS that we used as a basis.
@@
Towards closing the gap.  Compared to SemanticKITTI ($23.6$ PQ, $40\%$), we observe a significantly smaller gap between the model trained on pseudo- and ground truth labels for nuScenes ($31.3$ PQ, $44\%$). Despite its sparse Lidar signal and more diverse input scenarios, SAL performs better on this challenging Lidar dataset. We attribute this to a larger pseudo-label coverage of nuScenes ($48\%$, compared to $14\%$ coverage on SemanticKITTI), and, therefore, effectively more labeled data, which is also consistent with our observations on SemanticKITTI w.r.t. increasing the amount of labeled data (Tab.˜2). This trend suggests a clear path forward: Applying SAL to pseudo-label more data with zero annotation cost.
@@
Beyond labeled datasets.  Datasets, such as SemanticKITTI and nuScenes, provide dense, point-level panoptic segmentation labels, which we use for analysis and the evaluation of our method on a small set of classes that were labeled in the respective dataset ($14/11$ things/stuff in SemanticKITTI, respectively, and $23/6$ in nuScenes). To demonstrate the versatility of SAL, we additionally pseudo-label and train a model on Waymo Open [75] dataset and show qualitative results in Fig.˜1 and Appendix˜E. This effort takes less than a week of pseudo-labeling and model training efforts (details in Sec.˜B.2).
@@
The good, the (breaking) bad, and the ugly.  While we are thrilled about class-agnostic segmentation performance, our experimental analysis reveals a substantial gap between full supervision and our zero-shot model. Finally, at the moment, SAL is limited to training on specific datasets. As in fully-supervised Lidar perception, cross-sensor generalization remains a challenge for SAL. This could be addressed by utilizing temporal context, data scaling efforts and investigating sensor-agnostic backbone networks. We are excited about the new research opportunities that this work opens.
@@ 5 Conclusions
Acknowledgments.  This project was funded, in parts, by ERC Starting Grant DynAI (ERC-101043189). We are grateful to Žan Gojčič, Guillem Braso, Cristiano Saltori, Sérgio Agostinho, and Jonas Schult for their feedback on the paper and their insightful comments. Special thanks to Maxim Maximov for his help on figures!
@@
Appendix A Implementation Details
A.1 Pseudo-label Engine
@@
Mask generation.  We start by generating an overlapping set of segmentation masks for each image (camera view) (Algorithm˜1, L6) using Segment Anything (SAM) foundation model [31]. To generate masks, we utilize parameters, as described in Tab.˜A.1. This step yields $171$ masks per image on average. We then flatten SAM’s output mask hierarchy by non-maxima suppression (NMS) with a small overlap threshold (Algorithm˜1, L7, govern by NMS IoU threshold in Tab.˜A.1) to obtain a small, non-overlapping set of masks per-scan ($45$ on average after flatening). This way, we obtain a set of non-overlapping binary masks $m_{t}^{k}\in\{0,1\}^{W\times H}$ for each camera view $k$ with an image plane of size $W\times H$.
@@
In practice, during mask suppression, we sort masks based on the mask area (rather than score, usually done in NMS) – this criterion favors objects over their parts and subparts. As can be seen in Tab.˜B.3, this approach (NMS area) performs significantly better as compared to objectness score-based suppression (NMS score), which may favor object parts over objects. We visualize the flattened segmentation masks in images in Fig.˜E.3 (SemanticKITTI) and Fig.˜E.4 (nuScenes).
@@
CLIP image token generation.  We proceed by generating localized CLIP [67] image feature tokens $f_{t}^{k}\in\mathbb{R}^{768}$ for each binary mask $m_{t}^{k}$ (Algorithm˜1, L8). To obtain image tokens for a masked region of the input image, we utilize MaskCLIP [16] and their relative mask attention in the CLIP image encoder feature space. In Fig.˜E.3 and Fig.˜E.4, we visualize for each generated mask in the image the most-likely class, assigned using the generated CLIP token (according to the class vocabularies, as detailed in Tab.˜E.1). Importantly, we report class names only for visualization purposes in Fig.˜E.3, Fig.˜E.4, and Fig.˜E.5. We never classify segments in the image domain but merely distill CLIP image feature tokens to Lidar to facilitate zero-shot classification during inference where we do not use any image features.
@@
Image-to-Lidar unprojection.  From the Lidar perspective, we unproject each image mask $m_{t,i}^{k}\in\{0,1\}^{W\times H}$ to a binary Lidar segmentation mask $\tilde{m}^{k}_{t,i}\in\{0,1\}^{N}$ by transforming the respective camera coordinate frame to the Lidar space (Algorithm˜1, L9–L10). The unprojection yields pairs $\{\tilde{m}_{t}^{k},f_{t}^{k}\}$ of Lidar masks and their corresponding CLIP features w.r.t. camera $k$. On average, from $45$ masks generated in images, $39$ are transferred to the Lidar domain – the rest are either too small or not supported by Lidar measurements (due to signal sparsity or lack of Lidar coverage). We visualize masks, unprojected from a single camera to Lidar in the middle row of Fig.˜E.6 (SemanticKITTI).
@@
We then insert these masks to the output sets $\tilde{m}_{t},f_{t}$ as follows. To all masks in $\tilde{m}_{t}^{k}$ that do not overlap significantly overlap with masks in $\tilde{m}_{t}$ (Tab.˜A.1, multi-view IoU threshold), we assign a new ID, and insert masks and their corresponding CLIP features to the output sets $\tilde{m}_{t},f_{t}$. For masks whose overlap threshold exceeds this limit, we update existing masks with a union of the two, and the feature with the average. We visualize masks, unprojected to Lidar from several views, followed by fusion, in the middle row of Fig.˜E.7 (nuScenes).
@@
Refinement via clustering.  Finally, we improve our pseudo-label quality by creating an ensemble of DBSCAN [17] clusters $\tilde{m}_{t}^{DBSCAN}\in\{0,1\}^{N}$, obtained by varying the density thresholds (Tab.˜A.1, DBSCAN density thresholds) to compensate for varying density in Lidar point clouds. We replace each $\tilde{m}_{i}\in\tilde{m}_{t}$ with its best-matching $\tilde{m}_{l}\in\tilde{m}_{t}^{DBSCAN}$ in case their IoU exceeds a minimal overlap threshold (Tab.˜A.1, DBSCAN IoU overlap threshold) and retain the original mask otherwise to obtain a refined set of pseudo-labels that retain their original cardinality and associated CLIP features.
@@
DBSCAN clustering implementation.  To create the ensemble of DBSCAN clusters $\tilde{m}_{t}^{DBSCAN}$, we first perform geometric plane fitting and remove ground points (to estimate the ground points, we use [34] and its publicly-available implementation). We then perform DBSCAN on the ground-filtered point clouds using six density thresholds (reported in Tab.˜A.1, DBSCAN IoU overlap threshold). This leads to a large set of (overlapping) segmentation masks, induced by the estimated point clusters (in SemanticKITTI, $5,413$ on average per scan)
@@
In Tab.˜B.3, we report two alternatives. Firstly, rather than using DBSCAN, we perform erosion in the image domain to minimize the “edge bleeding” artifacts. As can be seen, this variant decreases the PQ score ($46.0$ to $42.2$). Second, rather than replacing segments with DBSCAN (DBSCAN replace), we filter the pool of SAM-generated segments, only retaining those that have a sufficient overlap (in terms of intersection-over-union), (DBSCAN filter). While this variant improves PQ for things classes, it significantly reduces PQ for stuff classes, leading to overall lower performance compared to the replace variant ($46.7$ PQ for filter vs. $48.7$ PQ for replace).
@@
Parameter tuning.  We tune all hyperparameters on a subset of $40$ Lidar scans of SemanticKITTI validation set (we sample every $100^{th}$ scan). We perform no parameter tuning for nuScenes dataset.
@@
6:  for each camera $k$ do
7:   $\mathcal{I}_{t,k}$ $\leftarrow$ load_image($t,k$)
8:   $m_{t}^{k}$ $\leftarrow$ SAM($\mathcal{I}_{t,k}$) // Generate masks in image $k$
9:   $m_{t}^{k}$ $\leftarrow$ NMS($m_{t}^{k}$) // Apply Non-Maximum Suppression (NMS) to masks
10:   $f_{t}^{k}\leftarrow$ MaskCLIP($\mathcal{I}_{t,k}$, $m_{t}^{k}$) // Obtain localized CLIP features for each mask
11:   $\tilde{m}_{t}^{k}\leftarrow$ label_point_cloud ($P_{t,k},m_{t}^{k}$, $c_{k}$) // Generate segmentation pseudo-labels w.r.t. camera $k$
12:   $\tilde{m}_{t},f_{t}\}\leftarrow$ insert_or_merge($\tilde{m}_{t},f_{t},\tilde{m}_{t}^{k}$, $f_{t}^{k}$) // Cross-camera fusion: Insert new or merge with existing
13:  end for
14:end for
[…]
A.2 Zero-shot Model
Table A.1: SAL hyperparameters. We show parameters for both components of our framework: (i) SAM model [31], which we use to generate segmentation masks in images; (ii) the pseudo-label generation engine and (iii) our zero-shot model. For the latter, we only highlight parameters that deviate from [45].
@@ A.2 Zero-shot Model
DBSCAN IoU overlap threshold $0.1$
@@
Figure A.1: SAL zero-shot model. Our model takes Lidar point clouds and text prompts as inputs. Its architecture relies on a 3D sparse-convolutional Minkowski backbone [12] followed by a Transformer decoder for object instance segmentation. The decoder computes cross-attention between object queries and backbone features. Three task heads predict objectness scores, segmentation masks, and CLIP tokens for each query. Once trained, we forward the dataset class vocabulary through the CLIP text encoder and perform zero-shot classification via matching with predicted CLIP tokens. The model requires no retraining for different vocabularies.
We use our SAL model, trained using labels generated by the pseudo-label engine (Sec.˜A.1), to segment and classify any object. The model implements a Transformer decoder architecture and adds a CLIP token prediction head (see Fig.˜A.1). In contrast to MaskPLS [45], our decoder operates not in point but voxel space and only on a single backbone feature scale. These adaptations obtained empirically better results from an overall more lightweight model. As it is common for Transformer decoders, the computation of its loss $\mathcal{L}_{\text{{{SAL}}{}}}$ relies on a bipartite cost matching between predictions and (pseudo-)labels. Empirically, we obtain better results if this matching only relies on the class-agnostic segmentation and does not incorporate the image token prediction as an additional cost. Unlike [45], we train all models from scratch and do not initialize our backbone with SegContrast [54]. In Tab.˜A.1, we highlight hyperparameter choices that deviate from the default setup in [45].
@@
A.3 Text Prompt Engineering
To perform zero-shot classification within a pre-defined class vocabulary, we complement and enrich the otherwise ambiguous and uninformative class names. As shown in Tab.˜E.1, each class is predictable not only by its own name but a set of additional prompts. In particular, all other-X classes are ambiguous prompts. The otherness only works in a fully supervised setting where a model can learn to predict, for example, all vehicles except the types already covered by other classes. In our case, we must directly prompt for other vehicle types, such as trailer, bus, tram, or train. This problem could also be solved by adding negations/exclusions to prompts. However, this approach did not yield the desired outcome in our experiments. We apply the same set of super classes for SemanticKITTI [6] and nuScenes [18] and merge their respective default class prompts. Furthermore, we follow [67] and wrap every prompt into a list of full-sentence templates. This results in text prompts like a photo of a car, which better align with the image caption training data of the CLIP text encoder. In Sec.˜C.1, we show additional ablations on the aforementioned text prompt engineering.

### v2

@@ 4.4 Lidar Panoptic Segmentation Evaluation
Finally, we compare SAL with state-of-the-art LPS methods on standard benchmarks [6, 18]. We focus this discussion on comparing SAL, trained using labeled data, to our SAL zero-shot model. For the zero-shot evaluation, we specify classes, defined in respective dataset vocabularies [6, 18], as text prompts (we detail these prompts in Sec.˜0.C.1). Importantly, zero-shot results are obtained without training on any labeled data. In contrast, all baselines are trained using human-labeled data and are constrained to their respective pre-defined class vocabularies.
@@
SemanticKITTI.  When evaluating with the default class vocabulary (Tab.˜5), our zero-shot model reaches $24.8$ PQ or $42\%$ of the fully-supervised model ($59.5$ PQ). As discussed in Sec.˜4.2, our method is mainly limited by the quality of the generated CLIP features. In practice, our zero-shot semantic classifier mainly confuses objects within super classes, e.g., car vs. other-vehicle. Therefore, without any retraining, we prompt our zero-shot model with super classes and obtain $48.5$ PQ, $59\%$ of the supervised model ($81.7$ PQ).
@@
nuScenes.  We report our result in Tab.˜5 and reach similar conclusions. SAL reaches $38.4$ PQ on default classes, $54\%$ of the supervised model ($70.5$ PQ). As with SemanticKITTI, results consistently improve when evaluated with super-classes. In this setting, zero-shot model yields $52.6$ PQ. Note that SAL model, supervised with labeled data, significantly outperforms MaskPLS.
@@
Towards closing the gap.  Compared to SemanticKITTI ($24.8$ PQ, $42\%$), we observe a significantly smaller gap between the model trained on pseudo- and ground truth labels for nuScenes ($38.4$ PQ, $54\%$). Despite its sparse Lidar signal and more diverse input scenarios, SAL performs better on this challenging Lidar dataset. We attribute this to a larger pseudo-label coverage of nuScenes ($48\%$, compared to $14\%$ coverage on SemanticKITTI), and, therefore, effectively more labeled data, which is also consistent with our observations on SemanticKITTI w.r.t. increasing the amount of labeled data (Tab.˜2). This trend suggests a clear path forward: applying SAL to pseudo-label more data with zero annotation cost.
@@
Beyond labeled datasets.  Datasets, such as SemanticKITTI and nuScenes, provide dense, point-level panoptic segmentation labels, which we use for analysis and the evaluation of our method on a small set of classes that were labeled in the respective dataset ($14/11$ things/stuff in SemanticKITTI, respectively, and $23/6$ in nuScenes). To demonstrate the versatility of SAL, we additionally pseudo-label and train a model on Waymo Open [75] dataset and show qualitative results in Fig.˜1 and Appendix˜0.E. This effort takes less than a week of pseudo-labeling and model training efforts (we provide details in Sec.˜0.B.2).
@@
The good, the (breaking) bad, and the ugly.  While we are thrilled about class-agnostic segmentation performance, our experimental analysis reveals a substantial gap between full supervision and our zero-shot model. Finally, at the moment, SAL is limited to training on specific datasets. As in fully-supervised Lidar perception, cross-sensor generalization remains a challenge for SAL. This could be addressed by utilizing temporal context, data scaling efforts and investigating sensor-agnostic backbone networks.
@@ 5 Conclusions
Acknowledgments.  This project was funded, in parts, by ERC Starting Grant DynAI (ERC-101043189). We are grateful to Žan Gojčič, Guillem Braso, Cristiano Saltori, Sérgio Agostinho, and Jonas Schult for their feedback on the paper and their insightful comments. Special thanks to Maxim Maximov for his help on figures.
@@
Appendix 0.A Implementation Details
0.A.1 Pseudo-label Engine
@@
Mask generation.  We start by generating an overlapping set of segmentation masks for each image (camera view) (Algorithm˜1, L6) using Segment Anything (SAM) foundation model [31]. To generate masks, we utilize parameters, as described in Tab.˜0.A.1. This step yields $171$ masks per image on average. We then flatten SAM’s output mask hierarchy by non-maxima suppression (NMS) with a small overlap threshold (Algorithm˜1, L7, govern by NMS IoU threshold in Tab.˜0.A.1) to obtain a small, non-overlapping set of masks per-scan ($45$ on average after flatening). This way, we obtain a set of non-overlapping binary masks $m_{t}^{k}\in\{0,1\}^{W\times H}$ for each camera view $k$ with an image plane of size $W\times H$.
@@
In practice, during mask suppression, we sort masks based on the mask area (rather than score, usually done in NMS) – this criterion favors objects over their parts and subparts. As can be seen in Tab.˜0.B.3, this approach (NMS area) performs significantly better as compared to objectness score-based suppression (NMS score), which may favor object parts over objects. We visualize the flattened segmentation masks in images in Fig.˜0.E.3 (SemanticKITTI) and Fig.˜0.E.4 (nuScenes).
@@
CLIP image token generation.  We proceed by generating localized CLIP [67] image feature tokens $f_{t}^{k}\in\mathbb{R}^{768}$ for each binary mask $m_{t}^{k}$ (Algorithm˜1, L8). To obtain image tokens for a masked region of the input image, we utilize MaskCLIP [16] and their relative mask attention in the CLIP image encoder feature space. The original MaskCLIP pipeline forwards the entire image and all masks at once. For our use case, we observed better per-mask classification results by generating image tokens for each mask separately. To this end, we forward individual image crops. In Fig.˜0.E.3 and Fig.˜0.E.4, we visualize for each generated mask in the image the most-likely class, assigned using the generated CLIP token (according to the class vocabularies, as detailed in Tab.˜0.E.1). Importantly, we report class names only for visualization purposes in Fig.˜0.E.3, Fig.˜0.E.4, and Fig.˜0.E.5. We never classify segments in the image domain but merely distill CLIP image feature tokens to Lidar to facilitate zero-shot classification during inference where we do not use any image features.
@@
Image-to-Lidar unprojection.  From the Lidar perspective, we unproject each image mask $m_{t,i}^{k}\in\{0,1\}^{W\times H}$ to a binary Lidar segmentation mask $\tilde{m}^{k}_{t,i}\in\{0,1\}^{N}$ by transforming the respective camera coordinate frame to the Lidar space (Algorithm˜1, L9–L10). The unprojection yields pairs $\{\tilde{m}_{t}^{k},f_{t}^{k}\}$ of Lidar masks and their corresponding CLIP features w.r.t. camera $k$. On average, from $45$ masks generated in images, $39$ are transferred to the Lidar domain – the rest are either too small or not supported by Lidar measurements (due to signal sparsity or lack of Lidar coverage). We visualize masks, unprojected from a single camera to Lidar in the middle row of Fig.˜0.E.6 (SemanticKITTI).
@@
We then insert these masks to the output sets $\tilde{m}_{t},f_{t}$ as follows. To all masks in $\tilde{m}_{t}^{k}$ that do not overlap significantly overlap with masks in $\tilde{m}_{t}$ (Tab.˜0.A.1, multi-view IoU threshold), we assign a new ID, and insert masks and their corresponding CLIP features to the output sets $\tilde{m}_{t},f_{t}$. For masks whose overlap threshold exceeds this limit, we update existing masks with a union of the two, and the feature with the average. We visualize masks, unprojected to Lidar from several views, followed by fusion, in the middle row of Fig.˜0.E.7 (nuScenes).
@@
Refinement via clustering.  Finally, we improve our pseudo-label quality by creating an ensemble of DBSCAN [17] clusters $\tilde{m}_{t}^{DBSCAN}\in\{0,1\}^{N}$, obtained by varying the density thresholds (Tab.˜0.A.1, DBSCAN density thresholds) to compensate for varying density in Lidar point clouds. We replace each $\tilde{m}_{i}\in\tilde{m}_{t}$ with its best-matching $\tilde{m}_{l}\in\tilde{m}_{t}^{DBSCAN}$ in case their IoU exceeds a minimal overlap threshold (Tab.˜0.A.1, DBSCAN IoU overlap threshold) and retain the original mask otherwise to obtain a refined set of pseudo-labels that retain their original cardinality and associated CLIP features.
@@
DBSCAN clustering implementation.  To create the ensemble of DBSCAN clusters $\tilde{m}_{t}^{DBSCAN}$, we first perform geometric plane fitting and remove ground points (to estimate the ground points, we use [34] and its publicly-available implementation). We then perform DBSCAN on the ground-filtered point clouds using six density thresholds (reported in Tab.˜0.A.1, DBSCAN IoU overlap threshold). This leads to a large set of (overlapping) segmentation masks, induced by the estimated point clusters (in SemanticKITTI, $5,413$ on average per scan)
@@
In Tab.˜0.B.3, we report two alternatives. Firstly, rather than using DBSCAN, we perform erosion in the image domain to minimize the “edge bleeding” artifacts. As can be seen, this variant decreases the PQ score ($46.0$ to $42.2$). Second, rather than replacing segments with DBSCAN (DBSCAN replace), we filter the pool of SAM-generated segments, only retaining those that have a sufficient overlap (in terms of intersection-over-union), (DBSCAN filter). While this variant improves PQ for things classes, it significantly reduces PQ for stuff classes, leading to overall lower performance compared to the replace variant ($46.7$ PQ for filter vs. $48.7$ PQ for replace).
@@
Parameter tuning.  We tune all hyperparameters on a subset of $40$ Lidar scans of SemanticKITTI validation set (we sample every $100^{th}$ scan). We perform no parameter tuning for nuScenes.
@@
6:  $\tilde{m}_{t}^{DBSCAN}$ $\leftarrow$ DBSCAN_ensamble($P_{t}$) // Generate (overlapping) Lidar mask ensemble
7:  for each camera $k$ do
8:   $\mathcal{I}_{t,k}$ $\leftarrow$ load_image($t,k$)
9:   $m_{t}^{k}$ $\leftarrow$ SAM($\mathcal{I}_{t,k}$) // Generate masks in image $k$
10:   $m_{t}^{k}$ $\leftarrow$ NMS($m_{t}^{k}$) // Apply Non-Maximum Suppression (NMS) to masks
11:   $f_{t}^{k}\leftarrow$ MaskCLIP($\mathcal{I}_{t,k}$, $m_{t}^{k}$) // Obtain localized CLIP features for each mask
12:   $\tilde{m}_{t}^{k}\leftarrow$ label_point_cloud ($P_{t,k},m_{t}^{k}$, $c_{k}$) // Generate segmentation pseudo-labels w.r.t. camera $k$
13:   $\tilde{m}_{t}^{k}\leftarrow$ replace($\tilde{m}_{t}^{k},\tilde{m}_{t}^{DBSCAN}$) // Replace with sufficiently overlapping DBSCAN masks
14:   $\tilde{m}_{t},f_{t}\}\leftarrow$ insert_or_merge($\tilde{m}_{t},f_{t},\tilde{m}_{t}^{k}$, $f_{t}^{k}$) // Cross-camera fusion: Insert new or merge with existing
15:  end for
16:end for
0.A.2 Zero-shot Model
Table 0.A.1: SAL hyperparameters. We show parameters for both components of our framework: (i) SAM model [31], which we use to generate segmentation masks in images; (ii) the pseudo-label generation engine and (iii) our zero-shot model. For the latter, we only highlight parameters that deviate from [45].
@@ 0.A.2 Zero-shot Model
DBSCAN IoU overlap threshold $0.5$
@@
Figure 0.A.1: SAL zero-shot model. Our model takes Lidar point clouds and text prompts as inputs. Its architecture relies on a 3D sparse-convolutional Minkowski backbone [12] followed by a Transformer decoder for object instance segmentation. The decoder computes cross-attention between object queries and backbone features. Three task heads predict objectness scores, segmentation masks, and CLIP tokens for each query. Once trained, we forward the dataset class vocabulary through the CLIP text encoder and perform zero-shot classification via matching with predicted CLIP tokens. The model requires no retraining for different vocabularies.
We use our SAL model, trained using labels generated by the pseudo-label engine (Sec.˜0.A.1), to segment and classify any object. The model implements a Transformer decoder architecture and adds a CLIP token prediction head (see Fig.˜0.A.1). In contrast to MaskPLS [45], our decoder operates not in point but voxel space and only on a single backbone feature scale. These adaptations obtained empirically better results from an overall more lightweight model. As it is common for Transformer decoders, the computation of its loss $\mathcal{L}_{\text{{{SAL}}{}}}$ relies on a bipartite cost matching between predictions and (pseudo-)labels. Empirically, we obtain better results if this matching only relies on the class-agnostic segmentation and does not incorporate the image token prediction as an additional cost. Unlike [45], we train all models from scratch and do not initialize our backbone with SegContrast [54]. In Tab.˜0.A.1, we highlight hyperparameter choices that deviate from the default setup in [45].
@@
0.A.3 Text Prompt Engineering
To perform zero-shot classification within a pre-defined class vocabulary, we complement and enrich the otherwise ambiguous and uninformative class names. As shown in Tab.˜0.E.1, each class is predictable not only by its own name but a set of additional prompts. In particular, all other-X classes are ambiguous prompts. The otherness only works in a fully supervised setting where a model can learn to predict, for example, all vehicles except the types already covered by other classes. In our case, we must directly prompt for other vehicle types, such as trailer, bus, tram, or train. This problem could also be solved by adding negations/exclusions to prompts. However, this approach did not yield the desired outcome in our experiments. We apply the same set of super classes for SemanticKITTI [6] and nuScenes [18] and merge their respective default class prompts. Furthermore, we follow [67] and wrap every prompt into a list of full-sentence templates. This results in text prompts like a photo of a car, which better align with the image caption training data of the CLIP text encoder. In Sec.˜0.C.1, we show additional ablations on the aforementioned text prompt engineering.